If the origin is transferred to the point (-3,2) keeping the axes parallel, then the equation 2x-3y+6=0 becomes-
A) 2x'+3y'=6
B) 2x'+3y'+6=0
C) 3x'-2y'=6
D) 2x'-3y'=6

Grade:11

1 Answers

Aditya Gupta
4 years ago
if origin is translated to (h,k) then x= X+h and y= Y+k.
so  2x-3y+6=0 can be written as 2( X – 3)+6= 3(Y+2)
or 2X= 3Y+6
or  2X – 3Y=6 
so option is correct.
kindly approve :)